OpenStudy (anonymous):

Felix exclaims that his quadratic with a discriminant of −1 has no real solutions. Felix then puts down his pencil and refuses to do any more work. Create an equation with a negative discriminant. Then explain to Felix, in calm and complete sentences, how to find the solutions, even though they are not real.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Have you created an equation with a negative discriminant ? i.e. where b^2 -4ac is < 0

OpenStudy (anonymous):

nope :l

OpenStudy (anonymous):

so 1^2 - 4(2)(3)?

OpenStudy (anonymous):

b = 1 a = 2 and c = 3?

OpenStudy (phi):

now write down that quadratic , and set = to zero

OpenStudy (anonymous):

how do I write it as a quadratic?

OpenStudy (phi):

the a,b and c are the coefficients of x^2, x and x^0

OpenStudy (anonymous):

ax^2 + bx + c is the quadratic formula right?

OpenStudy (anonymous):

b = 1 a = 2 and c = 3 2x^2 + 1x + 3

OpenStudy (phi):

no, that is the standard form of a quadratic the quadratic formula is how you solve it. *** now write down that quadratic , and set = to zero ***

OpenStudy (anonymous):

so what? 2x^2 + 1x + 3 = 0 would be it?

OpenStudy (phi):

it is understood that the "solutions" are what x values make the quadratic = 0 now that you have \[ 2x^2 + x + 3 = 0 \] (people don't bother to write a 1 as a coefficient... no number means 1) you solve it You could use "complete the square" to solve it.

OpenStudy (anonymous):

what do we do first in the complete the square method?

OpenStudy (anonymous):

ahh 14min lol

OpenStudy (phi):

It is a great video...especially if you have popcorn with it.

OpenStudy (anonymous):

ok I got past the part where he completed the square

OpenStudy (anonymous):

\[2x^2 + 1x + 3 = 0\] we already have 3 terms on the left side though o_O

OpenStudy (anonymous):

you there @phi?

OpenStudy (phi):

If you got far enough, you should divide both sides by 2 (to get a coefficient of 1 in front of x^2)

OpenStudy (anonymous):

so 1x^2 + 1x + 3 = 0

OpenStudy (anonymous):

or just x^2 + 1x + 3 = 0

OpenStudy (phi):

divide *everything* by 2

OpenStudy (anonymous):

ok x^2 + 1/2x + 3/2 = 0

OpenStudy (phi):

now try to finish it.

OpenStudy (anonymous):

x^2 + 2x = 0 -2x x^2 = -2?

OpenStudy (phi):

start with x^2 + 1/2x + 3/2 = 0 now review the video.

OpenStudy (anonymous):

we square 1/2 and 3/2?

OpenStudy (anonymous):

which is 1/4 and 9/4?